# Adjoint log canonical foliated singularities on surfaces

### v1

@@ 2.5. Log canonical foliated singularities
$\iota(E):=\begin{cases}1,&\text{if $E$ is not $\mathcal{F}$-invariant},\\ 0,&\text{if $E$ is $\mathcal{F}$-invariant}.\end{cases}$
@@ 2.6. $(\epsilon,\delta)$-adjoint log canonical foliated singularities
Next we introduce two important examples.
Example 2.15.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$ and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=ny\,{\rm d}x-x\,{\rm d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over the point $p$ (see Figure 1).
$\cdots\cdots$
$\bar{E}_{2}(-2)$
$\bar{E}_{1}(-2)$
$\bar{E}_{n-1}(-2)$
$E_{n}(-1)$
Figure 1.
We denote by $q_{i}$ (resp. $E_{i}$), $i=1,\cdots,n$, the successive blow-up points (resp. the exceptional curves). In this case, we have
In particular, each $E_{i}$ is $\mathcal{F}^{\prime}$-invariant for $i\leq n-1$ and $E_{n}$ is not $\mathcal{F}^{\prime}$-invariant. A direct computation gives
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{n},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+\cdots+(n-1)\bar{E}+nE_{n}.$
Therefore,
$K_{\mathcal{F}^{\prime}}+\epsilon K_{X^{\prime}}=\sigma^{*}(K_{\mathcal{F}}+K_{X})+\epsilon\sum_{i=1}^{n-1}i\bar{E}_{i}+(n\epsilon-1)E_{n}.$
For any $\epsilon\in(0,1)$, there exists some $n$ such that $n\epsilon-1\geq 0$. In this case, $p$ is $\epsilon$-adjoint canonical but not canonical.
Example 2.16.
Let $p$ be a singularity of a foliated surface $(X,\mathcal{F})$. Assume that $X$ is smooth at $p$ and that $\mathcal{F}$ is locally defined near $p=(0,0)$ by the $1$-form
$\omega=x\,\mathrm{d}x+y^{2}\,\mathrm{d}y.$
Consider the minimal resolution $\sigma\colon(X^{\prime},\mathcal{F}^{\prime})\to(X,\mathcal{F})$ over the point $p$ (see Figure 2).
$\bar{E}_{1}(-3)$
$\bar{E}_{2}(-2)$
$E_{3}(-1)$
Figure 2.
We denote by $q_{i}$ (resp. $E_{i}$) the successive blow-up points (resp. the exceptional curves). In this case, we have
$l(q_{1})=l(q_{2})=1\qquad\text{and}\qquad l(q_{3})=a(q_{3})=2.$
In particular, each $E_{i}$ is $\mathcal{F}^{\prime}$-invariant. A direct computation gives
$K_{\mathcal{F}^{\prime}}$ $=\sigma^{*}(K_{\mathcal{F}})-E_{3},$
$K_{X^{\prime}}$ $=\sigma^{*}(K_{X})+\bar{E}_{1}+2\bar{E}_{2}+4E_{3}.$
Therefore,
@@ 2.7. Notation and basic configurations
    ${\rm Z}(\mathcal{F},\Gamma_{1})=1$ and ${\rm Z}(\mathcal{F},\Gamma_{i})=0$ for all $i\geq 2$.
@@
An $\mathcal{F}$-dihedral graph is a divisor
@@
Figure 5. $\mathcal{F}$-dihedral graph
@@
Definition 2.23 ($\mathcal{F}$-star chain).
An $\mathcal{F}$-star chain is an $\mathcal{F}$-star graph whose underlying divisor is a chain.
@@ 3. Negative definite configurations of curves
In this section, we collect several standard facts on negative definite configurations of curves on smooth surfaces, which will be used repeatedly in the study of exceptional divisors arising in resolutions of singularities.
Throughout this section, let $\mathfrak{E}=\bigcup_{i=1}^{n}E_{i}$ be a finite collection of proper irreducible curves on a smooth surface $X$ whose intersection matrix $(E_{i}\cdot E_{j})$ is negative definite. Set $E:=E_{1}+\cdots+E_{n}$.
@@ 3.1. Dual graphs and intersection equations
Let $D$ be a divisor on $X$, and let $F=C_{1}+\cdots+C_{r}\subset E$ be a divisor whose intersection matrix is negative definite. Then the system of linear equations
@@
We denote by $M(D,F)$ the unique divisor supported on $F$ satisfying
@@ 3.2. $D^{>0}$-chains
and hence Lemma 2.25 implies $M(D,F+C)\geq M(D,F)$. Thus $\gamma^{\prime}_{i}>0$ for all $i=1,\dots,r+1$, and $F+C$ is a $D^{>0}$-chain.
@@
For later estimates along chains, we introduce the following determinants associated to $F$.
(3.1) $\begin{cases}\mu_{0}:=\lambda_{r+1}=1,\qquad\mu_{1}:=\lambda_{r}=1,\\ \mu_{k}:=\det(-C_{i}\cdot C_{j})_{1\leq i,j\leq k-1},&k=2,\dots,r+1,\\ \lambda_{l}:=\det(-C_{i}\cdot C_{j})_{l+1\leq i,j\leq r},&l=0,\dots,r-1,\\ n:=\lambda_{0}=\mu_{r+1}.\end{cases}$
Lemma 3.4.
  1. (1)
    For $i=1,\dots,r$, we have
    $\lambda_{i-1}-e_{i}\lambda_{i}+\lambda_{i+1}=0,\qquad\mu_{i-1}-e_{i}\mu_{i}+\mu_{i+1}=0,$
    where $e_{i}:=-C_{i}^{2}$.
  2. (2)
    For $i<j$, the quantity $\lambda_{i}\mu_{j}-\lambda_{j}\mu_{i}$ is a positive multiple of $n$. In particular,
    $\lambda_{i}\mu_{i+1}-\lambda_{i+1}\mu_{i}=n.$
Proof.
This follows from a straightforward computation. ∎
Since the intersection matrix $(C_{i}\cdot C_{j})_{1\leq i,j\leq r}$ is negative definite, for each $i$ there exists a unique effective $\mathbb{Q}$-divisor $M_{i}(F)$ supported on $F$ such that
$M_{i}(F)\cdot C_{i}=-1,\qquad M_{i}(F)\cdot C_{j}=0\quad(j\neq i).$
A direct computation yields
$M_{i}(F)=\frac{\lambda_{i}}{n}\sum_{k=1}^{i}\mu_{k}C_{k}+\frac{\mu_{i}}{n}\sum_{k=i+1}^{r}\lambda_{k}C_{k}.$
Consequently,
$M(D,F)=\sum_{i=1}^{r}(-D\cdot C_{i})\,M_{i}(F).$
Moreover, we have the following lemma:
Lemma 3.5.
Writing $M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i}$, we have
(3.2) $\gamma_{i}=\frac{\lambda_{i}}{n}\sum_{k=1}^{i}\mu_{k}\,(-D\cdot C_{k})+\frac{\mu_{i}}{n}\sum_{k=i+1}^{r}\lambda_{k}\,(-D\cdot C_{k}),\qquad i=1,\dots,r.$
Lemma 3.6.
Assume that $D\cdot C_{i}\geq 0$ for $i\geq 2$. Then $F$ is a $D^{>0}$-chain if and only if
$\sum_{k=1}^{r}\mu_{k}\,D\cdot C_{k}<0.$
In this case, for any $k\geq 2$,
$D\cdot C_{k}<\frac{-D\cdot C_{1}}{\mu_{k}}.$
Proof.
Using (3.2), we may write
$\gamma_{i}=\frac{\lambda_{i}}{n}\sum_{k=1}^{r}\mu_{k}\,(-D\cdot C_{k})-\sum_{k=i+1}^{r}\frac{\lambda_{i}\mu_{k}-\lambda_{k}\mu_{i}}{n}\,(-D\cdot C_{k}).$
By Lemma 3.4, we have $\lambda_{i}\mu_{k}-\lambda_{k}\mu_{i}\geq n$ for $k>i$. Since $D\cdot C_{k}\geq 0$ for $k\geq 2$, it follows that
$\gamma_{i}\geq\frac{\lambda_{i}}{n}\sum_{k=1}^{r}\mu_{k}\,(-D\cdot C_{k})=\lambda_{i}\gamma_{r}.$
Hence $F$ is a $D^{>0}$-chain if and only if $\gamma_{r}>0$, which is equivalent to $\sum_{k=1}^{r}\mu_{k}\,D\cdot C_{k}<0$.
In this case, for $i\geq 2$,
$\mu_{i}\,D\cdot C_{i}\leq\sum_{k=2}^{r}\mu_{k}\,D\cdot C_{k}<\mu_{i}\,(-D\cdot C_{1}),$
and the desired inequality follows. ∎
@@ 3.3. Special $D^{>0}$-chains
The next proposition provides a numerical criterion for positivity along special $D^{>0}$-chains, which will be crucial in the classification.
@@
Let $F=C_{1}+\cdots+C_{r}$ be a special $D^{>0}$-chain. Write $M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i}$. Then:
@@
    $\gamma_{k}\leq\frac{(-D\cdot C_{1})\lambda_{k}}{n},\qquad k=1,\dots,r.$
@@
    For each $k=1,\dots,r$, the sub-chain $F_{k}=C_{1}+\cdots+C_{k}$ is a special $D^{>0}$-chain, and
@@
    In particular,
    $D\cdot C_{k}<M(D,F_{k-1})\cdot C_{k}\leq\frac{-D\cdot C_{1}}{\mu_{k}},\qquad k\geq 2,$
    where $\mu_{k}=\det(-C_{i}\cdot C_{j})_{1\leq i,j\leq k-1}$.
@@
(1) follows directly from (3.2). (2) follows from Lemma 3.2 and Lemma 3.6. ∎
4. Classification of $\epsilon$-adjoint lc singularities with $\epsilon\in(0,1/3)$
Let $(Y,\mathcal{G})$ be a foliated pair on a normal surface $Y$ and let $p\in Y$ be a singular point of $(Y,\mathcal{G})$. Let
$\pi:(X,\mathcal{F})\longrightarrow(Y,\mathcal{G})$
be the minimal resolution of $(Y,\mathcal{G})$ at $p$. Denote by $\mathfrak{E}=\bigcup_{i\in I}E_{i}$ the set of $\pi$-exceptional curves and set
$E:=\sum_{i\in I}E_{i}.$
@@ 4. Classification of $\epsilon$-adjoint lc singularities with $\epsilon\in(0,1/3)$
$K_{(X,\mathcal{F},E),\epsilon}=\pi^{*}K_{(Y,\mathcal{G}),\epsilon}+\sum_{i}a_{i}E_{i}.$
Intersecting both sides with $E_{j}$, we obtain the system of linear equations
@@
It follows that $p$ is an $\epsilon$-adjoint log canonical (resp. $\epsilon$-adjoint klt) singularity if and only if
$a_{i}\geq 0\quad(\text{resp.\ }a_{i}>0)\quad\text{for all }i.$
Our goal is to classify all possible configurations of the exceptional divisor $E$ arising from $\epsilon$-adjoint lc (resp. klt) singularities. Equivalently, this amounts to determining all possible dual graphs of $E$ which are $K^{\geq 0}$-graphs (resp. $K^{>0}$-graphs), where

### v2

@@ 2.5. Log canonical foliated singularities
$\iota(E):=\begin{cases}1,&\text{if $E$ is not $\mathcal{F}^{\prime}$-invariant},\\ 0,&\text{if $E$ is $\mathcal{F}^{\prime}$-invariant}.\end{cases}$
@@ 2.7. Notation and basic configurations
    ${\rm Z}(\mathcal{F},\Gamma_{1})=1$ and ${\rm Z}(\mathcal{F},\Gamma_{i})=2$ for all $i\geq 2$.
@@
Lemma 2.17.
If $C$ is a (1)(-1)-$\mathcal{F}$-curve with $C^{2}=-1$, then $C$ is $\mathcal{F}$-exceptional. Moreover, there exists exactly one singularity of $\mathcal{F}$ on $C$, which is non-degenerate and reduced.
Lemma 2.18.
Suppose $C$ is a (2)(-2)-$\mathcal{F}$-curve with $C^{2}=-1$. Assume that every singularity of $\mathcal{F}$ on $C$ is reduced. If $C$ is not $\mathcal{F}$-exceptional, then at the contraction point $p$ of $C$, the induced foliation is locally defined by
$\omega=(x+y)\,{\rm d}x-x\,{\rm d}y.$
Moreover, there exists exactly one singularity of $\mathcal{F}$ on $C$, which is a saddle-node $q$. In a neighborhood of $q$, the foliation $\mathcal{F}$ is defined by
$\omega^{\prime}=y(s+1)\,{\rm d}s+s^{2}\,{\rm d}y,$
where $C=(y=0)$ is the weak separatrix of $q$.
Proof.
See [Bru15, pp. 7–8]. ∎
@@
An $\mathcal{F}$-dihedral fork is a divisor
@@
Figure 3. $\mathcal{F}$-dihedral fork
@@
Moreover, if $s\leq 2$, we call $F$ an $\mathcal{F}$-star chain; if $s\geq 3$, we call $F$ is an $\mathcal{F}$-star fork.
@@
2.8. Classification of lc surfaces singularities and lc foliated surface singularities
In this subsection, we recall the classification of log canonical surface singularities and log canonical foliated surface singularities, serving as a reference point for comparison with the $\epsilon$-adjoint setting.
Theorem 2.24.
Let $p$ be a log canonical singularity of a normal surface $Y$. Let
$\pi:X\to Y$
  1. (1)
    $E$ is a Hirzebruch-Jung chain.
  2. (2)
  3. (3)
    $E$ is a cycle of smooth rational curves.
  4. (4)
    $E$ has two forks with the dual graph as follows:
[…]
  5. (5)
    1. (a)
    2. (b)
      $l=3$ and $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, $(3,3,3)$.
    3. (c)
      $l=4$ and $(m_{1},m_{2},m_{3},m_{4})=(2,2,2,2)$.
Here, types (2), (3), (4), (5-b), and (5-c) are not klt, while types (1) and (5-1) are klt.
Proof.
See [Alex92] or [KM98, Sec. 4.1]. ∎
Theorem 2.25.
Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface with log canonical foliated singularity at $p$, and let $\pi:(X,\mathcal{F})\to(Y,\mathcal{G})$ be the minimal resolution with exceptional divisors $E=\cup_{i}E_{i}$. Then $E$ is one of the following:
  1. (1)
    An $\mathcal{F}$-chain. (See Figure 1.)
  2. (2)
    A chain of (2)(-2)-$\mathcal{F}$-curves. (See Definition 2.16.)
  3. (3)
  4. (4)
    An $\mathcal{F}$-dihedral fork: two (1)(-1)-$\mathcal{F}$-curves with self-intersection 2-2 joined by a bad tail connecting to a chain of (2)(-2)-$\mathcal{F}$-curves. (See Figure 3.)
  5. (5)
  6. (6)
Here, type (1) is terminal and types (1)–(5) are canonical (as foliated singularities).
Proof.
See [Chen23, Theorem 1.1]. We note that cases (6) and (7) in [Chen23, Theorem 1.1] are combined into a single case, namely case (6) in Theorem 2.25. ∎
@@ 3. Negative definite configurations of curves
In this section, we introduce useful notations and properties for negative definite configurations of curves on a smooth surface, which will be applied in the next section.
Throughout this section, let $\mathfrak{E}=\bigcup_{i=1}^{n}E_{i}$ be a finite collection of irreducible curves on a smooth surface $X$ whose intersection matrix $(E_{i}\cdot E_{j})$ is negative definite. Set $E:=E_{1}+\cdots+E_{n}$.
@@ 3.1. Dual graphs and intersection equations
Let $D$ be a divisor on $X$, and let $F=C_{1}+\cdots+C_{r}\subset E$ be a divisor whose intersection matrix is negative definite. Since the intersection matrix is invertible, the system of linear equations
@@
The divisor $M(D,F)$ is the unique divisor supported on $F$ satisfying
@@
The divisor $M(D,F)$ can be viewed as the intersection-theoretic projection of $D$ onto the negative definite subconfiguration $F$.
@@
Lemma 3.2.
Assume the negative definite curves $F=\sum_{i=1}^{r}E_{i}$ is connected. If $DE_{i}\leq 0$ for all $i$, then $F$ is a $D^{\geq 0}$-graph. Moreover, $F$ is a $D^{>0}$-graph if and only if $DE_{i}<0$ for some $i$.
Proof.
See [KM98, Lemma 3.41]. ∎
@@ 3.2. $D^{>0}$-chains
and hence Lemma 2.26 implies $M(D,F+C)>M(D,F)$; in particular, $\gamma^{\prime}_{r+1}>0$. Thus $\gamma^{\prime}_{i}>0$ for all $i=1,\dots,r+1$, and $F+C$ is a $D^{>0}$-chain.
@@
For later numerical control along chains, we introduce the determinants
(3.1) $\begin{cases}\mu_{0}=\lambda_{r+1}:=1,\qquad\mu_{1}=\lambda_{r}:=1,\\ \mu_{k}:=\det(-C_{i}\cdot C_{j})_{1\leq i,j\leq k-1},&k=2,\dots,r+1,\\ \lambda_{l}:=\det(-C_{i}\cdot C_{j})_{l+1\leq i,j\leq r},&l=0,\dots,r-1,\\ n:=\lambda_{0}=\mu_{r+1}.\end{cases}$
@@ 3.3. Special $D^{>0}$-chains
We now isolate the configurations that play a decisive role in the later classification.
@@
Here $C_{1}$ is called the first curve of $F$.
The following proposition provides explicit and uniform bounds for the coefficients of $M(D,F)$ along special $D^{>0}$-chains. More precisely, each coefficient is controlled solely by the initial negativity $D\cdot C_{1}$ and the combinatorial structure of the chain.
@@
Let $F=C_{1}+\cdots+C_{r}$ be a special $D^{>0}$-chain. Write
$M(D,F)=\sum_{i=1}^{r}\gamma_{i}C_{i}.$
Then:
@@
    For $k=1,\dots,r$, we have
    $\gamma_{k}\leq\frac{(-D\cdot C_{1})\lambda_{k}}{n}.$
@@
    Let $F_{k}:=C_{1}+\cdots+C_{k}$ for $k=1,\dots,r$. Then for $k\geq 2$,
    $D\cdot C_{k}<M(D,F_{k-1})\cdot C_{k}\leq\frac{-D\cdot C_{1}}{\mu_{k}}.$
    Moreover, each $F_{k}$ is a special $D^{>0}$-chain, satisfying
@@
(1) Let $M_{1}(F)$ be the unique effective $\mathbb{Q}$-divisor supported on $F$ such that
$M_{1}(F)\cdot C_{1}=-1,\qquad M_{1}(F)\cdot C_{i}=0\ \text{ for }i\geq 2.$
A straightforward computation shows that
$M_{1}(F)=\sum_{k=1}^{r}\frac{\lambda_{k}}{n}C_{k}.$
On the other hand, we have
$\big[(-D\cdot C_{1})M_{1}(F)-M(D,F)\big]\cdot C_{i}\leq 0,\qquad i=1,\dots,r.$
By Lemma 2.26, this implies
$M(D,F)\leq(-D\cdot C_{1})M_{1}(F)=\sum_{k=1}^{r}\frac{(-D\cdot C_{1})\lambda_{k}}{n}C_{k}.$
This proves (1).
(2) For $k\leq r-1$, we compute
$\big[M(D,F)-M(D,F_{k})\big]\cdot C_{i}=\begin{cases}0,&\text{if }i\leq k,\\[4.0pt] D\cdot C_{k+1}-M(D,F_{k})\cdot C_{k+1},&\text{if }i=k+1,\\[4.0pt] D\cdot C_{i}\ (\geq 0),&\text{if }i\geq k+2.\end{cases}$
If
$D\cdot C_{k+1}-M(D,F_{k})\cdot C_{k+1}\geq 0,$
then Lemma 2.26 implies $M(D,F)\leq M(D,F_{k})$, hence $\gamma_{i}\leq 0$ for all $i\geq k+1$, contradicting the assumption that $F$ is a $D^{>0}$-chain. Therefore,
(3.2) $D\cdot C_{k}<M(D,F_{k-1})\cdot C_{k}\leq\frac{-D\cdot C_{1}}{\mu_{k}},\qquad k\geq 2,$
where the second inequality follows from the same argument as in (1).
Note that $F_{1}$ is a special $D^{>0}$-chain since $D\cdot C_{1}<0$. By (3.2) and Lemma 3.3, an induction on $k$ shows that each $F_{k}$ is a special $D^{>0}$-chain and that $M(D,F_{k})$ is strictly increasing. ∎
4. Classification of $\epsilon$-adjoint log canonical singularities for $\epsilon\in(0,1/3)$
In the previous section, we analyzed negative definite configurations of curves on smooth surfaces and established several structural properties of their intersection matrices. We now apply these results to the study of $\epsilon$-adjoint log canonical surface singularities.
Let $(Y,\mathcal{G})$ be a foliated pair on a normal surface, and let $p$ be a singular point of $(Y,\mathcal{G})$. Fix $\epsilon\in(0,1/3)$. Assume that $(Y,\mathcal{G},p)$ is $\epsilon$-adjoint log canonical. Our goal is to determine all possible configurations arising in its minimal resolution.
Let $\pi:X\to Y$ be the minimal resolution of the germ $(Y,\mathcal{G},p)$ (cf. Definition 2.5), and write $\mathfrak{E}=\bigcup_{i\in I}E_{i}$ for the exceptional divisor. Set
$E=\sum_{i\in I}E_{i}.$
Note that the intersection matrix $(E_{i}\cdot E_{j})_{i,j\in I}$ is negative definite. Consequently, the adjoint log canonical condition can be expressed as a system of linear inequalities for the discrepancies along the components $E_{i}$.
@@ 4. Classification of $\epsilon$-adjoint log canonical singularities for $\epsilon\in(0,1/3)$
$K_{(X,\mathcal{F},E),\epsilon}=\pi^{*}K_{(Y,\mathcal{G}),\epsilon}+\sum_{i\in I}a_{i}E_{i},$
where $K_{(X,\mathcal{F},E),\epsilon}$ and $K_{(Y,\mathcal{G}),\epsilon}$ are as in Definition 2.13. Intersecting both sides with $E_{j}$, we obtain
@@
Since the intersection matrix is negative definite, the system (4.1) admits a unique solution $(a_{i})$. Hence $(Y,\mathcal{G},p)$ is $\epsilon$-adjoint log canonical (resp. $\epsilon$-adjoint klt) if and only if
$a_{i}\geq 0\quad(\text{resp.\ }a_{i}>0)\quad\text{for all }i\in I.$
Thus the classification problem reduces to determining all negative definite configurations for which the right-hand side of (4.1) produces a nonnegative (resp. positive) solution. Equivalently, in the terminology of Section 3, we must determine all possible dual graphs of $E$ which are $K^{\geq 0}$-graphs (resp. $K^{>0}$-graphs), where
@@
The remainder of this section is devoted to a systematic analysis of these possibilities.